OpenStudy (anonymous):

miguel is earning extra money by painting houses. he charges a $200 fee plus $12 per can of paint needed to complete the job. write and use an equation to find how many cans of paint he needs for a $260 job

OpenStudy (anonymous):

200+12c=260 12c=260-200 12c=60 c=60/12 c=5